A grasshopper leaps into the air from the edge of a vertical cliff. The values for the parameters are d = 6.33cm, theta = 56.0 deg and R = 1.82m

A. What is the initial speed of the grass hopper
B. What is the height of the cliff h

Please give height of cliff in m

Edit R Notes itio 12-17 AM Handwritten note Notes: 8/27/17 We should an access point ot thet. Current Problems-.-.
Answer #1

H2= Ho2 + 2g*h = 0
Ho2 - 19.6*0.0633 = 0
Ho2 = 1.24
Ho = 1.11 m/s = Ver. component of initial velocity.

Ho = Vo*sin560 = 1.11
Vo = 1.11/sin560 = 1.34 m/s = Initial velocity.

Xo = Vo*Cos56 = 1.11*Cos56 = 0.622 m/s =
Hor. component of initial velocity.

Dx = Xo*T = 1.82 m.
0.622*T = 1.82
T = 2.92 s. = Fall time.

h = 0.5g*T2 - 0.0633 = 4.9*2.922 - 0.0633 = 41.88 m. = 42m =Ht. of the cliff.